# ./references.bib

### v1

@@ Abstract
A variety of physical phenomena, such as amplification, absorption, and radiation, can be effectively described using non-Hermitian operators. However, the introduction of non-uniform non-Hermiticity can lead to the formation of exceptional points in a system’s spectrum, where two or more eigenvalues become degenerate and their associated eigenvectors coallesce causing the underlying operator or matrix to become defective. Here, we explore extensions of the Clifford and quadratic $\epsilon$-pseudospectrum, previously defined for Hermitian operators, to accommodate non-Hermitian operators and matrices, including the possibility that the underlying operators may possess exceptional points in their spectra. In particular, we provide a framework for finding approximate joint eigenstates of a $d$-tuple of Hermitian operators $\bm{A}$ and non-Hermitian operators $\bm{B}$, and show that their Clifford and quadratic $\epsilon$-pseudospectra are still well-defined despite any non-normality. We prove that the non-Hermitian quadratic gap is local with respect to the probe location when there are perturbations to one or more of the underlying operators. Altogether, this framework enables the exploration of non-Hermitian physical systems’ $\epsilon$-pseudospectra, including but not limited to photonic systems where gain, loss, and radiation are prominent physical phenomena.
@@ 1 Introduction
Localized states in physical systems have a variety of important technological applications. In multidimensional systems, states that are localized to a single dimension are useful for enabling directed transport of energy or information. Moreover, zero-dimensional cavity states have uses in storing energy and information. In systems described by Hermitian Hamiltonians, finding such localized states is generally accomplished by identifying (or designing) one or more states to be isolated in the system’s spectrum, whose corresponding local density of states (LDOS) is spatially localized. Indeed, finding classes of systems that exhibit such boundary-localized states is one of the primary motivations for studying material topology [bansil_lin_das_RevModPhys.88.021004],[hasan_kane_2010_RevModPhys.82.3045],[ozawa_et_all_RevModPhys.91.015006],[xiao_chang_niu_RevModPhys.82.1959], where bulk-boundary correspondence can guarantee such localized states.
Open systems have the property that the corresponding Hamiltonian is non-Hermitian, stemming from either a loss of energy or decoherence due to the environment. As such, for an open system, the Hamiltonian $H$ can have complex valued eigenvalues whose imaginary parts correspond to the loss (or gain) of energy or information in each associate eigenstate. A canonical class of open systems are photonic systems, which can exhibit amplification, absorption, and radiation. Here, without including microscopic descriptions of the sources of amplification or absorption, or considering the infinite space into which photons can radiate, these phenomena are instead generally considered by introducing non-Hermitian terms into the photonic system’s Hamiltonian. Nevertheless, finding low-loss localized states in photonics, and open systems more broadly, is a perennial goal, as these states exhibit Purcell enhancement of any latent nonlinear effects [purcell_spontaneous_1946].
Previously, the search for system states that are approximately localized in energy and position, i.e., that are approximate joint eigenstates of both a system’s Hamiltonian and position operators, has been studied using the system’s quadratic pseudospectrum [cerjan_loring_vides_2023]. In such an approach, the idea is to choose a $(d+1)$-tuple $\bm{\lambda}=(x_{1},...,x_{d},E)$ (or by abuse of notation also written as $(\bm{x},E)$) consisting of position and energy and extracting information from the quadratic composite operator
@@
The smallest eigenvalue will then provide a unit state that is an approximate joint eigenstate of each of the matrices $X_{i}$ and $H$. In particular the ‘probe site’ $\bm{\lambda}$ in $\mathbb{R}^{n+1}$, provides the corresponding approximate eigenvalues of the matrices. This mathematical method gives rise to the quadratic $\epsilon$-pseudospectrum and is the main tool in [cerjan_loring_vides_2023]. Similar approaches based on the system’s Clifford $\epsilon$-pseudospectrum have also been used to both identify localized states in realistic photonic systems and develop a class of invariants for material topology [cerjan_loring_pho_2022, kahlil_cerjan_loring, cerjan_prl_2024]. By itself, the inclusion of non-Hermiticity does not preclude the possibility of finding localized states generally, nor boundary localized states protected by material topology see [Esaki_Sato_Hasebe_Kohmoto_PR_2011], [kawabata_shiozaki_ueda_sato_PhysRevX.9.041015_2019], [Kawabata_etal_nature_2019], and [Shen_Zhen_Fu_PRL_2018]. However, when working with a non-Hermitian systems, we can fail to find an approximately localized state when directly applying the method outlined above with the $\epsilon$-pseudospectrum. The problem is that the quadratic composite operator in Equation 1 can fail to be normal for non-Hermitian Hamiltonians $H$ and as a result could have degenerate eigenspaces. Thus the degeneracy in the eigenspaces and eigenvalues of Equation 1 could imply the quadratic gap is not small enough. Thus we could fail to find approximately localized states since the error bounds guaranteeing a good approximately localized eigenstate depends on the magnitude of the quadratic gap function.
@@
Pseudospectral methods are being used for non-Hermitian systems, see [kiorpelidis2024scaling]. Here, we propose pseudospectral methods for predicting approximately localized states and resonances in non-Hermitian systems in both position and energy. We modify the quadratic composite operator in [cerjan_loring_vides_2023] to a generalized version of the following,
$\text{Q}_{(\bm{x},E)}\left(\bm{X},H\right)=\sum_{i}^{d}(X_{i}-x_{i})^{2}+(H-E)^{\dagger}(H-E)$ (2)
@@
while allowing $E$ to be complex valued since $H$ being non-Hermitian gives us the possibility of complex eigenvalues. This updates definitions for a probe site as well as the $\epsilon$-pseudospectrum. We take this idea and expand it to allow for multiple non-Hermitian operators in the construction, in particular non-normal operators ($AA^{\dagger}\not=A^{\dagger}A$). We prove Lipschitz continuity of the corresponding gap functions. We also prove that under perturbations to a physical system’s Hamiltonian, the quadratic gap function allows us to find approximately localized states with small loss based on the probe site. Along the way, we investigate the Clifford composite operator and argue why the radial gap can be especially helpful in non-Hermitian systems with exceptional points. We also make a connection between the different gap functions and highlight the theory with numerical simulations. The theory and numerical simulations allow us to consider a system with an exceptional point and how the radial gap function helps spot the exceptional point. We also use the theory for numerical simulations on a line gapped system. Overall, one of our goals is to use the non-Hermitian pseudospectral methods we develop to design systems where we have some level of control over these long-lived lossy states. In particular, we anticipate that a modified quadratic $\epsilon$-pseudospectrum will have applications in designing photonic structures for enhancing light-matter interactions.
@@
The remainder of the paper is outlined as follows: In Section 2, we introduce the idea of joint eigenstates between two matrices or observables and review known results for when all matrices involved are Hermitian. We finalize by highlighting some ideas from [cerjan_loring_vides_2023]. In Section 3, we consider a specific construction of the non-Hermitian spectral localizer in [cerjan_koekebier_Schulz_Baldes], which we call the Clifford composite operator $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$, and note some salient properties. We also briefly review the non-Hermitian Clifford linear gap and introduce a new gap called the non-Hermitian Clifford radial gap; the idea being that the Clifford linear gap will be more useful for line gapped physical systems, while the clifford radial gap could be more useful for point gapped systems. Additionally, we prove some results about the gaps and finalize with a conjecture. In Section 4, we motivate the construct the quadratic composite operator $\text{Q}_{(\bm{\lambda},\nu)}\left(\bm{A},\bm{B}\right)$ in Equation 2 by computing
@@
We define a quadratic gap $\mu_{(\bm{\lambda},\nu)}^{\textup{Q}}\left(\bm{A},B\right)$ and prove it’s equivalence to other numerical values. We end the section by proving Lipschitz continuity of the quadratic gap as well as provide important corollaries about joint approximate eigenstates. In Section 5, we provide error estimates between the non-Hermitian Clifford linear gap, non-Hermitian Clifford radial gap, and quadratic gap. In Section 6, we show that the quadratic gap is local, meaning that far away perturbations of $B$ relative to probe site $(\bm{\lambda},\bm{\nu})$ will have little effect on the corresponding non-Hermitian quadratic gap at that point. In Section 7, we apply the non-Hermitian Clifford linear gap, radial gap, and non-Hermitian quadratic gap functions to a two level non-Hermitian System. We argue why the non-Hermitian radial gap function should be used instead of the gap mentioned in [cerjan_koekebier_Schulz_Baldes] in the case of point gapped systems. In Section 8, we take a similar tight binding model as the one used in [cerjan_koekebier_Schulz_Baldes], and numerically compute the Clifford linear and quadratic gap functions to visualize were we might find approximately localized states in position and energy. Finally, we compare the Clifford linear gap to the quadratic gap function which highlights how closely the functions agree and where.
2 Pseudospectrum for joint approximate eigenstates
@@
The main idea of pseudospectra and it’s techniques is to give equal treatment to each operator in a physical system. Instead of computing eigenstates for a Hamiltonian and then computing expectation values, pseudospectral techniques combine all of the incompatible observables/matrices in the system into a single composite operator. The spectrum of said composite operator give us information about states approximately localized at some position and energy when we consider a system’s position and Hamiltonian operators, see [krejcirik_siegl_tater_viola_2015_10.1063/1.4934378],[loring_schulz-baldes_2017_12504413820170101],[loring_schulz-baldes_2020_edsgcl.62678504020200301],[loring],[trefethen_1997_edsjsr.213303719970901],[trefethen_embree_2005_MR2155029],[trefethen_reddy_driscoll_1993_6dce06aa-3a1d-396a-ad07-b01838717a11]. When the matrices commute it can be show that there is a unit state that is an eigenstate (with potentially different eigenvalue) for each observable. This is a standard result in the theory of linear algebra. However it is also known that incompatible observables do not commmute, this tells us that there does not exist a joint eigenstate of the corresponding observables. The main idea however is that we are going to develop a framework in which we can find a unit eigenstate that is an approximate joint eigenstate of multiple matrices/observables through pseudospectral methods. In order to motivate the study of approximate joint eigenstates and as a result the study of $\epsilon$-pseudospectra; we begin with the study of joint eigenstates from commuting matrices/observables.
@@
In the study of linear algebra it is known that a $d$ tuple of diagonalizable matrices $\bm{A}$ pairwise commute if and only if there exist a joint eigenstate. What this means is that if we have two matrices $A,B$, then the commutator is zero ($[A,B]=0$) then there is a state $\bm{\psi}$ such that $A\phi=\lambda_{A}\bm{\psi}$ and $B\bm{\psi}=\lambda_{B}\bm{\psi}$. To be more precise Theorem 2.3.3 in [horn_johnson] tells us that there is a unitary matrix $U$ that simultaneously transforms $A_{i}$ into upper triangular matrices, given a $d$ tuple of commuting normal matrices $\bm{A}$. In the situation were each matrix in $\bm{A}$ is normal ($A_{i}A_{i}^{\dagger}=A_{i}^{\dagger}A_{i}$) the theorem tell us that $A_{i}$ can be simultaneously diagonalized by the same unitary transformation.
@@
If $H$ is some Hamiltonian and $X$ is a position observable, we know that in general $[X,H]\not=0$. In the special case of an atomic limit, we do have $[X,H]=0$. However, most materials are not in the atomic limit. The situation of commutators of observables not commuting ($[X,H]\not=0$) is due to a fundamental idea that can be generalized to multiple dimensions. In particular $\left[\frac{d}{dx},x\right]\not=0$, in other words as operators the derivative and position don’t commute. The physics comes into play when we note that a Hamiltonian is usually a differential operator.
Thus if we have a physical system the best we can hope for are states that are approximately localized in position and energy. In other words if we have some number of position observables $\bm{X}$, and corresponding Hermitian Hamiltonian $H$, Proposition II.1 in [cerjan_loring_vides_2023] answers the question of looking for $(\bm{x},E)$, and $\bm{\psi}$ so that
@@
Indeed the approach that is taken is to utilize pseudospectral methods. Broadly speaking pseudospectra is a mathematical tool to determine approximate spectrum for non-commuting matrices. To find approximate joint eigenstates of matrices or observables, the idea is to consider various eigenvalue problems $(A_{i}-\lambda_{i})\bm{\psi}$, with $A_{i}$ Hermitian. Then we take a non trivial Clifford representation $\Gamma_{i}^{\dagger}=\Gamma_{i}$, $\Gamma_{i}^{2}=I$, $\Gamma_{i}\Gamma_{k}=-\Gamma_{k}\Gamma_{i}$ for $i\not=k$ to construct the so called spectral localizer mentioned in Section I of [cerjan_loring_vides_2023] as well as Section 1 of [loring]. Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}\in\mathbb{R}^{d+1}$,
@@
By properties of the tensor product and the $A_{i}$ matrices, the spectral localizer is itself a Hermitian matrix. From here a so called spectral localizer gap is defined and has alternative characterizations. It is defined explicitly in Section I of [cerjan_loring_vides_2023]. However it first appears implicitly in Definition 1.1 of a so called Clifford $\epsilon$-pseudospectrum in [loring].
@@
$L_{(\kappa x,\kappa y,\kappa z,E)}^{\textup{3d}}\left(X,Y,Z,H\right)=\kappa(X-x)\otimes\Gamma_{1}+\kappa(Y-y)\otimes\Gamma_{2}+\kappa(Z-z)\otimes\Gamma_{3}+(H-E)\otimes\Gamma_{4}.$
@@
for example. In any case it is good to note that there are other possible representations. This is proven in [cerjan_loring_2024_127892] by Cerjan and Loring.
@@
The smallest singular value of the spectral localizer plays a major role in an attempt at finding states that are localized in it’s various obeservables/operators. For this reason, in this context it is given a name. Specifically given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the spectral localizer gap is defined as
@@
This gap function is historically used implicity in the definition of the Clifford $\epsilon$-pseudospectrum that is used in Section 1 of [loring]. Then as it’s signifiance grew it warranted a definition, and was defined explicitly in Section I of [cerjan_loring_vides_2023]. In particular, given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators the Clifford $\epsilon$-pseudospectrum is defined as
@@
With all this in place, one of the main results of [loring] is finding a joint approximate states with corresponding probe site each corresponding to an element in the $\epsilon$-pseudospectrum. Lemma 1.2 and 1.3 in [loring] give formal error estimates, on the quantities $\left\lVert A_{i}\bm{\psi}-\lambda_{k}\bm{\psi}\right\rVert$.
We will prove a slightly generalized version of these two Lemmas in Section 3. The main idea being that a small gap, or the pseudospectrum provides us with a recipe to get our hands on a unit state $\bm{\psi}$ that is approximately localized in position and space.
Now suppose that the operators we are working with are in $\mathcal{M}_{n}(\mathbb{C})$, for every two more operators you have in your system the dimension of the spectral localizer doubles. For 2 and 3 observables we have the spectral localizer is in $\mathcal{M}_{2n}(\mathbb{C})$, for 4 and 5 observables the spectral localizer is in $\mathcal{M}_{4n}(\mathbb{C})$, etc. This could be an issue when it comes to numerical computations as $n$ can be very large to begin with. However, the details depend on the problem at hand.
The paper [cerjan_loring_vides_2023] tackles this by considering the fact that for Hermitian matrices $A$, $A\bm{\psi}=0$ iff $A^{2}\bm{\psi}=0$. Indeed if we square the spectral localizer we get
$\left(L_{\bm{\lambda}}(\bm{A})\right)^{2}=\left(\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})^{2}\otimes I\right)+\left(\sum_{i\not=k}[A_{i},A_{k}]\otimes\Gamma_{i}\Gamma_{k}\right).$
Indeed when the commutator terms are small in spectral norm, the predominant part of the squared Clifford composite operator is the first sum. This is a block diagonal matrix
@@
The eigenvalues of Equation 4 are determined by the repeated block matrix in 5.

### v2

# Clifford and quadratic composite operators with applications to non-Hermitian physics

@@ Abstract
A variety of physical phenomena, such as amplification, absorption, and radiation, can be effectively described using non-Hermitian operators. However, the introduction of non-uniform non-Hermiticity can lead to the formation of exceptional points in a system’s spectrum, where two or more eigenvalues become degenerate and their associated eigenvectors coalesce causing the underlying operator or matrix to become defective. Here, we explore extensions of the Clifford and quadratic $\epsilon$-pseudospectrum, previously defined for Hermitian operators, to accommodate non-Hermitian operators and matrices, including the possibility that the underlying operators may possess exceptional points in their spectra. In particular, we provide a framework for finding approximate joint eigenvectors of a $d$-tuple of Hermitian operators $\bm{A}$ and non-Hermitian operators $\bm{B}$, and show that their Clifford and quadratic $\epsilon$-pseudospectra are still well-defined despite any non-normality. We prove that the non-Hermitian quadratic gap is local with respect to the probe location when there are perturbations to one or more of the underlying operators. Altogether, this framework enables the exploration of non-Hermitian physical systems’ $\epsilon$-pseudospectra, including but not limited to photonic systems where gain, loss, and radiation are prominent physical phenomena.
@@ 1 Introduction
Localized states in physical systems have a variety of important technological applications. In multidimensional systems, states that are localized to a single dimension are useful for enabling directed transport of energy or information. Moreover, zero-dimensional cavity states have uses in storing energy and information. In systems described by Hermitian Hamiltonians, finding such localized states is generally accomplished by identifying (or designing) one or more states to be isolated in the system’s spectrum, whose corresponding local density of states (LDOS) is spatially localized. Indeed, finding classes of systems that exhibit such boundary-localized states is one of the primary motivations for studying material topology (See [2],[14],[25],[38]), where bulk-boundary correspondence can guarantee such localized states.
Open systems have the property that the corresponding Hamiltonian is non-Hermitian, stemming from either a loss of energy or decoherence due to the environment. As such, for an open system, the Hamiltonian $H$ can have complex eigenvalues whose imaginary parts correspond to the loss (or gain) of energy or information in each associated eigenvector. A canonical class of open systems are photonic systems, which can exhibit amplification, absorption, and radiation. Here, without including microscopic descriptions of the sources of amplification or absorption, or considering the infinite space into which photons can radiate, these phenomena are instead generally considered by introducing non-Hermitian terms into the photonic system’s Hamiltonian. Nevertheless, finding low-loss localized states in photonics, and open systems more broadly, is a perennial goal, as these states exhibit Purcell enhancement of any latent nonlinear effects [26].
Previously, the search for system states that are approximately localized in energy and position, i.e., that are approximate joint eigenvectors of both a system’s Hamiltonian and position operators, has been studied using the system’s quadratic pseudospectrum [4]. In such an approach, the idea is to choose a $(d+1)$-tuple $\bm{\lambda}=(x_{1},...,x_{d},E)=(\bm{x},E)$ consisting of position and energy and extracting information from the quadratic composite operator
@@
The smallest eigenvalue corresponds to a unit state that is an approximate joint eigenvector of each of the matrices $X_{i}$ and $H$. In particular, the ‘probe site’ $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, provides the corresponding approximate eigenvalues of the matrices. This mathematical method gives rise to the quadratic $\epsilon$-pseudospectrum and is the main tool in [4]. Similar approaches based on the system’s Clifford $\epsilon$-pseudospectrum have also been used to both identify localized states in realistic photonic systems and develop a class of invariants for material topology [6, 9, 5]. By itself, the inclusion of non-Hermiticity does not preclude the possibility of finding localized states generally, nor boundary localized states protected by material topology see [11], [17], [18], and [31]. However, when working with non-Hermitian systems, we can fail to find an approximately localized state when directly applying the method outlined above with the $\epsilon$-pseudospectrum. The problem is that the quadratic composite operator in Equation 1 can fail to be normal for non-Hermitian Hamiltonians $H$, and consequently could have degenerate eigenspaces. Thus, the degeneracy of the eigenspaces and eigenvalues of Equation 1 could imply the quadratic gap is not small enough. Thus we could fail to find approximately localized states, since the error bounds guaranteeing a good approximately localized eigenvector depend on the magnitude of the quadratic gap function.
@@
Pseudospectral methods are being used for non-Hermitian systems, see [19]. Here, we propose pseudospectral methods for predicting approximately localized states and resonances in non-Hermitian systems in both position and energy. We modify the quadratic composite operator in [4] to a generalized version of the following,
$Q_{(\bm{x},E)}\left(\bm{X},H\right)=\sum_{i}^{d}(X_{i}-x_{i})^{2}+(H-E)^{\dagger}(H-E)$ (2)
@@
while allowing $E$ to be complex valued since $H$ being non-Hermitian gives us the possibility of complex eigenvalues. This updates definitions for a probe site as well as the $\epsilon$-pseudospectrum. We take this idea and expand it to allow for multiple non-Hermitian operators in the construction, in particular non-normal operators ($AA^{\dagger}\not=A^{\dagger}A$). We prove Lipschitz continuity of the corresponding gap functions. We also prove that under perturbations to a physical system’s Hamiltonian, the quadratic gap function allows us to find approximately localized states with small loss based on the probe site. Along the way, we investigate the non-Hermitian spectral localizer and argue why the radial gap can be especially helpful in non-Hermitian systems with exceptional points. We also make a connection between the different gap functions and highlight the theory with numerical simulations. The theory and numerical simulations allow us to consider a system with an exceptional point and how the radial gap function helps spot the exceptional point. We also use the theory for numerical simulations on a line gapped system. Overall, one of our goals is to use the non-Hermitian pseudospectral methods we develop to design systems where we have some level of control over these long-lived lossy states. In particular, we anticipate that a modified quadratic $\epsilon$-pseudospectrum will have applications in designing photonic structures for enhancing light-matter interactions.
@@
The remainder of the paper is outlined as follows: In Section 2, we introduce the idea of joint eigenvectors between two matrices or observables and review known results for when all matrices involved are Hermitian. We finalize by highlighting some ideas from [4]. In Section 3, we consider a specific construction of the non-Hermitian spectral localizer in [3], which we call the non-Hermitian spectral localizer $L_{(\bm{\lambda},\nu)}\left(\bm{A},B\right)$, and note some salient properties. We also briefly review the non-Hermitian Clifford linear gap and introduce a new gap called the non-Hermitian Clifford radial gap; the idea being that the Clifford linear gap will be more useful for line gapped physical systems, while the clifford radial gap could be more useful for point gapped systems. Additionally, we prove some results about the gaps and finalize with a conjecture. In Section 4, we motivate the construction of the quadratic composite operator $Q_{(\bm{\lambda},\nu)}\left(\bm{A},\bm{B}\right)$ in Equation 2 by computing
@@
We define a quadratic gap $\mu_{(\bm{\lambda},\nu)}^{\textup{Q}}\left(\bm{A},B\right)$ and prove its equivalence to other numerical values. We end the section by proving Lipschitz continuity of the quadratic gap as well as provide important corollaries about joint approximate eigenvectors. In Section 5, we provide error estimates between the non-Hermitian Clifford linear gap, non-Hermitian Clifford radial gap, and quadratic gap. In Section 6, we show that the quadratic gap is local, meaning that far away perturbations of $B$ relative to probe site $(\bm{\lambda},\bm{\nu})$ will have little effect on the corresponding non-Hermitian quadratic gap at that point. In Section 7, we apply the non-Hermitian Clifford linear gap, radial gap, and non-Hermitian quadratic gap functions to a two level non-Hermitian system. We argue why the non-Hermitian radial gap function should be used instead of the gap mentioned in [3] in the case of point gapped systems. In Section 8, we take a similar tight binding model as the one used in [3], and numerically compute the Clifford linear and quadratic gap functions to visualize were we might find approximately localized states in position and energy. Finally, we compare the Clifford linear gap to the quadratic gap function which highlights how closely the functions agree and where.
2 Pseudospectrum for joint approximate eigenvectors
The use of pseudospectral methods is not a new one. Over the years it has been used either implicitly explicity. Some early uses of these methods with an eye towards physics is mentioned in [1],[21], and [22]. The ideas of pseudospectra became well known in the 90’s after Trefethen and Embree’s book [34]. This was the culmination of previous papers written by Trefethen, see [36],[33]. It is important to note that we only lightly touch on the history of the pseudospectrum and more can be read in [34]. When we consider the eigenvalue problem. In other words it is equivalent to asking if $(A-\lambda)^{-1}$ exists given a square matrix $A$ and some $\lambda\in\mathbb{C}$. However for numerical purposes it is better to ask if $\lVert(A-\lambda)^{-1}\rVert$ is large. This is how Trefethen and Embree motivate the pseudospectra of the matrix $A$ in [34]. In particular the proposed definition is that for some $\epsilon>0$ the pseudospectrum of $A$ is the set of complex numbers $\lambda$ such that $\lVert(A-\lambda)^{-1}\rVert>\epsilon^{-1}$ for some small $\epsilon>0$. In their book [34], Trefethen and Embree give alternative definitions that are all equivalent for matrices. Here we include them to for completeness.
Definition 2.1.
Suppose $A$ in $M_{n}(\mathbb{C})$ and $\epsilon>0$. The $\epsilon$-pseudospectrum denoted $\text{Spec}_{\epsilon}(A)$ of $A$ is the set of $z$ in $\mathbb{C}$ that satisfy $\lVert(A-z)^{-1}\rVert>\epsilon^{-1}$.
Definition 2.2.
Given $A$ in $M_{n}(\mathbb{C})$ the set $\text{Spec}_{\epsilon}(A)$ is the set of all $z$ in $\mathbb{C}$ such that $z$ in $\text{Spec}(A+E)$ for $E$ in $M_{n}(\mathbb{C})$ with $\lVert E\rVert<\epsilon$.
Definition 2.3.
Suppose $A$ in $M_{n}(\mathbb{C})$, $\text{Spec}_{\epsilon}(A)$ is the set of all $z$ in $\mathbb{C}$ such that $\lVert(A-z)\bm{v}\rVert<\epsilon$ for some non-zero $\bm{v}$ satisfying $\lVert\bm{v}\rVert=1$
Definition 2.4.
Suppose $A$ in $M_{n}(\mathbb{C})$, $\text{Spec}_{\epsilon}(A)$ is the set of all $z$ in $\mathbb{C}$ such that $\sigma_{\min}(A-z)<\epsilon$.
It is important to note that for definitions 2.1,2.2,2.3 can use matrix norms other than the operator norm. However all four definitions (Definition 2.1,2.2,2.3,2.4) are only equivalent when using the operator norm on matrices. Definition 2.3 and 2.4 in their book show up in an analogous way in the present paper. Moving beyond the works of Trefethen, the topic of joint spectra has been further investigated by Jhanjee in their PhD Thesis, see [16]. They investigate joint spectra for both matrices and operators in the commuting case. In the present paper we emphasize what happens in the non-commuting case for the matrix case.
@@ 2 Pseudospectrum for joint approximate eigenvectors
The main technique of pseudospectra in the multi matrix setting is to give equal treatment to each operator in a physical system. Instead of computing eigenvectors for a Hamiltonian and then computing expectation values, pseudospectral techniques combine all of the incompatible observables/matrices in the system into a single composite operator. The spectrum of said composite operator gives us information about states approximately localized at some position and energy when we consider a system’s position and Hamiltonian operators, see [20],[23],[24],[32],[33],[34],[35]. When the matrices commute it can be shown that there is a unit state that is an eigenvector (with potentially different eigenvalue) for each observable. This is a standard result in the theory of linear algebra. However, it is also known that incompatible observables do not commute; in other words we are not guaranteed of the existence of joint eigenvectors of the corresponding observables. The main idea however is that we are going to develop a framework in which we can find a unit eigenvector that is an approximate joint eigenvector of multiple matrices/observables through pseudospectral methods. In order to motivate the study of approximate joint eigenvectors and as a result the study of $\epsilon$-pseudospectra; we begin with the study of joint eigenvectors from commuting matrices/observables.
@@
It is known that a $d$ tuple of matrices $\bm{A}$ pairwise commute if and only if there there is a unitary that simultaneously transforms the $d$ matrices to upper traingular by a single unitary $U$ see Theorem 2.3.3 in [29]. This gives us the existence of at least one simultaneous eigenvector. Hence, if we have two matrices $A,B$ and the commutator is zero ($[A,B]=0$) then there is a state $\bm{\psi}$ such that $A\bm{\psi}=\lambda_{A}\bm{\psi}$ and $B\bm{\psi}=\lambda_{B}\bm{\psi}$. In the situation where each matrix in $\bm{A}$ is normal ($A_{i}A_{i}^{\dagger}=A_{i}^{\dagger}A_{i}$) the theorem tell us that $A_{i}$ can be simultaneously diagonalized by the same unitary transformation.
@@
If $H$ is some Hamiltonian and $X$ is a position observable, we know that in general $X,H]\not=0$. In the special case of an atomic limit, we have $[X,H]=0$. However, most materials are not in the atomic limit. We have $X,H]\not=0$ because $\left[\frac{d}{dx},x\right]\not=0$. The physics comes into play when we note that a Hamiltonian is usually a differential operator. Thus if we have a physical system the best we can hope for are states that are approximately localized in position and energy. In other words if we have some number of position observables $\bm{X}$, and corresponding Hermitian Hamiltonian $H$, Proposition II.1 in [4] answers the question of looking for $(\bm{x},E)$, and $\bm{\psi}$ so that
@@
Indeed the approach that is taken is to utilize pseudospectral methods. Broadly speaking pseudospectra is a mathematical tool to determine approximate spectrum for non-commuting matrices. To find approximate joint eigenvectors of matrices or observables, the idea is to consider various eigenvalue problems $(A_{i}-\lambda_{i})\bm{\psi}$, with $A_{i}$ Hermitian. Then we take a non trivial Clifford representation $\Gamma_{i}^{\dagger}=\Gamma_{i}$, $\Gamma_{i}^{2}=I$, $\Gamma_{i}\Gamma_{k}=-\Gamma_{k}\Gamma_{i}$ for $i\not=k$ to construct the so called spectral localizer mentioned in Section I of [4] as well as Section 1 of [32]. Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}\in\mathbb{R}^{d+1}$,
@@
By properties of the tensor product and the $A_{i}$ matrices, the spectral localizer is itself a Hermitian matrix. From here a so called spectral localizer gap is defined and has alternative characterizations. It is defined explicitly in Section I of [4]. However it first appears implicitly in Definition 1.1 of a so called Clifford $\epsilon$-pseudospectrum in [32].
@@
$L_{(\kappa x,\kappa y,\kappa z,E)}^{\textup{3d}}\left(\kappa X,\kappa Y,\kappa Z,\kappa H\right)=\kappa(X-x)\otimes\Gamma_{1}+\kappa(Y-y)\otimes\Gamma_{2}+\kappa(Z-z)\otimes\Gamma_{3}+(H-E)\otimes\Gamma_{4}.$
@@
for example. This is proven in [7] by Cerjan and Loring.
@@
The smallest singular value of the spectral localizer plays a major role in an attempt at finding states that are localized in its various observables/operators. For this reason, in this context it is given a name. Specifically given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators, and probe site $\bm{\lambda}$ in $\mathbb{R}^{d+1}$, the spectral localizer gap is defined as
@@
Given $\bm{A}$ in $\Pi_{i=1}^{d+1}\mathcal{M}_{n}(\mathbb{C})$ with each $A_{i}$ being Hermitian operators the Clifford $\epsilon$-pseudospectrum is defined as
@@
With all this in place, one of the main results of [32] is finding a joint approximate state with corresponding probe site each corresponding to an element in the $\epsilon$-pseudospectrum. Lemma 1.2 and 1.3 in [32] give formal error estimates, on the quantities $\left\lVert A_{i}\bm{\psi}-\lambda_{k}\bm{\psi}\right\rVert$. We prove a generalized version of these two lemmas in Section 3. The main idea being that a small gap or the pseudospectrum provides a recipe to get our hands on a unit state $\bm{\psi}$ that is approximately localized in position and space.
Now suppose that the operators in question are in $\mathcal{M}_{n}(\mathbb{C})$, The dimension of the spectral localizer is $n2^{\lfloor\frac{d+1}{2}\rfloor}\times n2^{\lfloor\frac{d+1}{2}\rfloor}$. Then for every two more matrices you have in your system the dimension of the spectral localizer becomes $n2^{\lfloor\frac{d+1}{2}\rfloor+1}\times n2^{\lfloor\frac{d+1}{2}\rfloor+1}$. For two and three observables the spectral localizer is in $\mathcal{M}_{2n}(\mathbb{C})$, for 4 and 5 observables the spectral localizer is in $\mathcal{M}_{4n}(\mathbb{C})$, etc. We can reduce the computational requirements by considering the fact that for hermitian matrices $A$, that $A\bm{\psi}=0$ iff $A^{2}\bm{\psi}=0$. Indeed if we square the spectral localizer we get
$\left(L_{\bm{\lambda}}(\bm{A})\right)^{2}=\left(\sum_{i=1}^{d+1}(A_{i}-\lambda_{i})^{2}\otimes I\right)+\left(\sum_{i<k}[A_{i},A_{k}]\otimes\Gamma_{i}\Gamma_{k}\right).$
When the commutator terms are small in operator norm, the predominant part of the squared non-Hermitian spectral localizer is the first sum. This is a block diagonal matrix
@@
The eigenvalues of Equation (4) are determined by the repeated block matrix in Equation (5).